# Linear Attention Sequence Parallelism

### v1

@@ 2.2 LASP
Backward Pass.
For the backward pass, let us examine the inverse procedure. We first define $\mathbf{dkv}_{t}$ for subsequent analysis. Given $\mathbf{do}_{s}$, we have (Katharopoulos et al., 2020)
$\mathbf{dq}_{s}^{\top}$ $=\mathbf{do}_{s}^{\top}\mathbf{kv}_{s}^{\top}\in\mathbb{R}^{1\times d},$ (13)
$\mathbf{dk}_{s}^{\top}$ $=\mathbf{v}_{s}^{\top}\mathbf{dkv}_{s}^{\top}\in\mathbb{R}^{1\times d},$
$\mathbf{dv}_{s}^{\top}$ $=\mathbf{k}_{s}^{\top}\mathbf{dkv}_{s}\in\mathbb{R}^{1\times d},$
$\mathbf{dkv}_{s}$ $=\sum_{i\geq s}\lambda^{i-s}\mathbf{q}_{i}\mathbf{do}_{i}^{\top}\in\mathbb{R}^{d\times d}.$
By writing $\mathbf{dkv}_{s}$ in a recursive form, we have
$\mathbf{dkv}_{n+1}$ $=0\in\mathbb{R}^{d\times d},$ (14)
$\quad\mathbf{dkv}_{s-1}$ $=\lambda\mathbf{dkv}_{s}+\mathbf{q}_{s-1}\mathbf{do}_{s-1}^{\top}.$
In the sequence parallelism scenario, we have $\{\mathbf{Q_{t}},\mathbf{K_{t}},\mathbf{V_{t}},\mathbf{O_{t}},\mathbf{dO_{t}}\}$ which corresponds to the $t$-th sub-sequence chunk on rank $i$, where $t\in\{1,\cdots,T\}$ and $i\in\{1,\cdots,W\}$. Same with the forward pass, the following derivations assume $t=i$, $T=W$.
$\small\mathbf{dQ}_{t,\mathrm{intra}}=[(\mathbf{dO}_{t}\mathbf{V}_{t}^{\top})\odot\mathbf{M}]\mathbf{K}_{t}.$ (15)
$\mathbf{dq}_{s+C}^{\top}=\mathbf{do}_{s+C}^{\top}\sum_{i\leq s+C}\lambda^{s+C-i}\mathbf{v}_{i}\mathbf{k}_{i}^{\top}$ (16)
$=$ $\mathbf{do}_{s+C}^{\top}\sum_{i=C+1}^{C+s}\lambda^{s+C-i}\mathbf{v}_{i}\mathbf{k}_{i}^{\top}+\lambda^{s}\mathbf{do}_{s+C}^{\top}\sum_{i\leq C}\lambda^{C-i}\mathbf{v}_{i}\mathbf{k}_{i}^{\top}.$
The first part (before the plus sign) in Eq. (16) corresponds to the intra-chunk, while the second part (after the plus sign) corresponds to the inter-chunk. In sequence parallelism scenario, we can calculate $\mathbf{dQ}_{t,\mathrm{inter}}$ as
$\small\mathbf{dQ}_{t,\mathrm{inter}}=\mathbf{\Lambda}\mathbf{dO}_{t}\mathbf{KV}_{t-1}^{\top}.$ (17)
Next, let us take $\mathbf{dK}$ into consideration, $\mathbf{dK}$ within the $t$-th chunk can be calculated in parallel as
$\small\mathbf{dK}_{t,\mathrm{intra}}=[(\mathbf{dO}_{t}\mathbf{V}_{t}^{\top})\odot\mathbf{M}]^{\top}\mathbf{Q}_{t}.$ (18)
Then we transform Eq. (13) as
$\mathbf{dk}_{s}^{\top}=\mathbf{v}_{s}^{\top}\sum_{i\geq s}\lambda^{i-s}\mathbf{do}_{i}\mathbf{q}_{i}^{\top}$ (19)
$=$ $\mathbf{v}_{s}^{\top}\sum_{i=s}^{C}\lambda^{i-s}\mathbf{do}_{i}\mathbf{q}_{i}^{\top}+\lambda^{C-s}\mathbf{v}_{s}^{\top}\sum_{i\geq C+1}\lambda^{i-C}\mathbf{do}_{i}\mathbf{q}_{i}^{\top},$
where the first term (before the plus sign) corresponds to the intra-chunk, and the second term (after the plus sign) corresponds to the inter-chunk. The above equation can be rewritten in terms of chunks as follow:
$\small\mathbf{dK}_{t,\mathrm{inter}}=\lambda^{C}\mathbf{\Lambda}^{-1}\mathbf{V}_{t}\mathbf{dKV}_{t+1}^{\top}.$ (20)
$\mathbf{dKV}_{t}$ $=\sum_{s>tC}\lambda^{s-tC}\mathbf{q}_{s}\mathbf{do}_{s}^{\top}$ (21)
$=\lambda^{C}\!\!\!\!\sum_{s>(t+1)C}\!\!\!\!\lambda^{s-(t+1)C}\mathbf{q}_{s}^{\top}\mathbf{do}_{s}+\!\!\!\!\sum_{s=tC+1}^{(t+1)C}\!\!\!\!\lambda^{s-tC}\mathbf{q}_{s}\mathbf{do}_{s}^{\top}$
Then a Send operation is performed to push $\mathbf{dKV}_{t}$ to rank $(i-1)$.
Finally, for $\mathbf{dV}$, its intra part can be calculated as
$\small\mathbf{dV}_{t,\mathrm{intra}}=[(\mathbf{Q}_{t}\mathbf{K}_{t}^{\top})\odot\mathbf{M}]^{\top}\mathbf{dO}_{t}.$ (22)
Again we transform Eq. (13) as:
$\mathbf{dv}_{s}^{\top}=\mathbf{k}_{s}^{\top}\sum_{i\geq s}\lambda^{i-s}\mathbf{q}_{i}\mathbf{do}_{i}^{\top}$ (23)
$=$ $\mathbf{k}_{s}^{\top}\sum_{i=s}^{C}\lambda^{i-s}\mathbf{q}_{i}\mathbf{do}_{i}^{\top}+\lambda^{C-s}\mathbf{k}_{s}^{\top}\sum_{i\geq C+1}\lambda^{i-C}\mathbf{q}_{i}\mathbf{do}_{i}^{\top}.$
The first and second terms (before and after the plus sign) corresponds to the computation of the intra- and inter-chunks, respectively. In sequence parallelism, $\mathbf{dV}_{t,\mathrm{inter}}$ can be calculated as:
$\small\mathbf{dV}_{t,\mathrm{inter}}=\lambda^{C}\mathbf{\Lambda}^{-1}\mathbf{K}_{t}\mathbf{dKV}_{t+1}.$ (24)
$\mathbf{dQ}_{t}$ $=\mathbf{dQ}_{t,\mathrm{intra}}+\mathbf{dQ}_{t,\mathrm{inter}},$ (25)
We provide the comprehensive formulation of the backward pass for LASP with $W=T$ in Algorithm 3.
2.3 Communication Analysis
Table 1: Communication Volume Comparison. Last column: we eliminate the common factors $Bd$ for ease of comparison.
Method Full Formulation Simplified Formulation
LASP $Bd^{2}/h$ $d/h$
DeepSpeed-Ulysses $4BNd/T$ $4N/T$
Megatron-SP $2BNd+4BNd/T$ $2N+4N/T$
When examining the LASP algorithm, it is important to note that the forward pass requires communication for the $\mathbf{KV}$ activation in each linear attention module layer. The communication volume is determined by $Bd^{2}/h$, where $B$ is the batch size and $h$ is the number of heads. In comparison, sequence parallelism in Megatron-LM utilizes all-gather operations twice after two layer normalization layers within each transformer layer, and a reduce-scatter operation after the attention and Feedforward Neural Network (FFN) layers. This results in a communication volume of $2BNd+4BNd/T$. DeepSpeed uses all-to-all collective communication operation (Thakur et al., 2005) for input $\mathbf{Q},\mathbf{K},\mathbf{V}$, and output $\mathbf{O}$ of each attention module layer, resulting in a communication volume of $4BNd/T$.
Table 1 displays a comparison of communication volumes across three frameworks. $d/h$ is the head dimension which is set at 128 as usual  (Lan et al., 2020). In practical applications where $N/T\geq 32$, LASP is able to achieve the lowest theoretical communication volume. Furthermore, the communication volume of LASP is not impacted by changes in sequence length $N$ or sub-sequence length $C$, which is a huge advantage for extremely long sequence parallelism across large GPU clusters.
2.4 System Engineering Optimization
$\mathbf{KV}$ State Caching. To avoid recomputing activation $\mathbf{KV}$ during the backward pass, we choose to store it in the HBM of the GPU right after computing it in the forward pass. During the subsequent backward pass, LASP directly accesses $\mathbf{KV}$ for use. It is important to note that the size of the $\mathbf{KV}$ activation cached in HBM is $d\times d$, which is not affected by the sequence length $N$. When the input sequence length $N$ is exceptionally large, the memory usage of $\mathbf{KV}$ becomes insignificant.
2.5 Data-Sequence Hybrid Parallelism
The technique of Data Parallelism is commonly used to split input data along the batch dimension for large-scale distributed deep learning. However, LASP offers a different approach by dividing data along the sequence dimension, which makes it easier to integrate with Data Parallelism. As explained in Section 2.2 and illustrated in Fig. 2, LASP allows for the specification of a smaller sequence parallel size that is divisible by the distributed world size. This configuration results in the input data being split along both the batch and sequence dimensions, which is a type of hybrid parallelism called data-sequence hybrid parallelism.
Emerging as significant distributed training techniques, sharded data parallelism methodologies aim to mitigate GPU memory usage during the training of large models. The ZeRO-series optimizers (Rajbhandari et al., 2020) in DeepSpeed and FSDP (Zhao et al., 2023) in PyTorch propose to distribute model states, which include optimizer states, gradients, and model parameters, across all GPUs within the distributed environment. This strategic distribution significantly reduces memory utilization on individual GPUs. As variants of data parallelism, these techniques seamlessly align with LASP. However, their focus on minimizing the memory of model states complements LASP’s objective of reducing activation memory on each GPU. By combining these methods, training large models with long sequence lengths becomes more feasible.
3 Related Work
Linear Attention. Linear Transformer models bypass the use of Softmax attention by adopting various approximation methods  (Katharopoulos et al., 2020; Choromanski et al., 2020; Peng et al., 2021; Qin et al., 2022b, a; Yang et al., 2023; Qin & Zhong, 2023; Qin et al., 2024c) instead. The central concept involves using the "kernel trick" to speed up the calculation of the attention matrix, specifically by multiplying keys and values before tackling the computationally intensive $n\times n$ matrix multiplication. For instance,  Katharopoulos et al. (2020) use $1+\mathrm{elu}$ activation function,  Qin et al. (2022b) utilizes the cosine function to imitate Softmax characteristics, and  Choromanski et al. (2020); Zheng et al. (2022, 2023) leverage sampling techniques to closely replicate the Softmax process are all strategies employed to achieve this.
Sequence Parallelism. Sequence parallelism is a widely used method to train long sequences for neural process language (NLP) tasks. This technique has been integrated into many large model training frameworks, including Megatron-LM, DeepSpeed, and Colossal-AI. Megatron-LM implements SP along with model (tensor) parallelism (MP) to perform large matrix multiplications on GPUs. However, MP partitions the attention heads, which limits the maximum parallelism degree to be less than the number of attention heads. DeepSpeed-Ulysses uses an all-to-all communication primitive to reduce communication volume, but also partitions attention heads and faces similar issues as Megatron-LM. Colossal-AI integrates ring-style communication called Ring Self-Attention (RSA) in SP to efficiently compute attention scores across devices. Recent LightSeq  (Li et al., 2023) introduces SP with load balancing and communication overlap scheduling, and a gradient checkpointing strategy to reduce activation memory further.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Scalability Evaluation of LASP on Throughput (tokens/sec) and Memory Usage. Left: Integration of LASP with FSDP backend; Right: Integration of LASP with DDP backend. The TNL-1B model is used, with a batch size of 1 across up to 128 A100 80GB GPUs. The sign "$\times$" with a dotted line represents occurring an Out of Memory (OOM).
4 Experiments
We evaluate LASP on two representative linear attention-based models: TransNormerLLM (TNL) (Qin et al., 2024a) and Linear Transformer (Katharopoulos et al., 2020). TNL is the latest large language model purely built upon linear attention, while Linear Transformer is a classical linear transformer model recognized in the community. Our assessment focuses on three key areas: I) the ability of LASP to scale up sequence length on scaling-out GPUs, II) the convergence when using LASP, and III) speed evaluation when using LASP and its comparison with other SP methods. No Activation Checkpointing (AC)  (Korthikanti et al., 2022) techniques are used in following experiments to reduce activation memory, except experiments in Section 4.3. This is because although the adoption of AC will further enables longer sequence lengths, it will cover up the ability of our sequence parallel method LASP. All experiments are conducted on a GPU cluster equipped with 128 A100 80G GPUs. Our implementation is built on MetaSeq (Zhang et al., 2022), a PyTorch-based sequence modeling framework with FairScale (FairScale authors, 2021) integrated. For more details of hardware and software, see Appendix A.1.
Experimental Setup.
@@ 4 Experiments
The training configuration is set with specific hyperparameters: a learning rate of 0.0005 to control the step size, a cap of 50,000 updates to define the training duration, and a 2,000-update warmup period to stabilize early training by gradually adjusting the learning rate (Zhou et al., 2020). Additionally, a weight decay rate of 0.01 is used for regularization to avoid over-fitting. The Adam optimizer, with beta values of 0.9 and 0.999, is chosen for managing the momentum and scaling of gradients, aiding in effective and stable training convergence. Different DDP backends, including PyTorch DDP (abbr. DDP), Legacy DDP, FSDP, ZeRO-series, are selected in experiments for cross-validation of compatibility with LASP.
Figure 4: Speed Comparison (tokens/sec) of LASP Against DeepSpeed-Ulysses and Megatron-SP. The sign "$\times$" with a dotted line represents occurring an Out of Memory (OOM). The evaluation utilizes the TNL-1B and 7B models with a batch size of 1 on 64 A100 80GB GPUs. The parallelism size for these three methods is configured to 64.
4.1 Scalability and Speed Comparison
The scalability results regarding throughput and memory usage with varying sequence lengths and number of GPUs are illustrated in Fig. 3. By using LASP, we successfully scale the sequence length up to 4096K using the FSDP backend and 2048K with the DDP backend on a TNL model with 1B parameters, on 128 GPUs.
Table 2: Convergence Performance of LASP. All experiments use 8 A100 80G GPUs, 16K sequence length, and batch size of 1. The results cover various DDP backends in conjunction with LASP. We explore the performance of two linear attention models: TransNormerLLM (TNL) and Linear Transformer, both with 0.4B parameters, across 50K updates.

### v3

@@ A.1 Backward Pass Algorithm
A.2 Hardware and Software
Hardware.
Our experimental configuration involves a maximum of 16x DGX-A100 servers, each equipped with 8x A100 GPUs, these GPUs are interconnected through NVSwitch, ensuring an inter-GPU bandwidth of 600GBps. For inter-node communication, we employ RoCE (RDMA over Converged Ethernet) technology, utilizing 8 RoCE RDMA adapters in each server. This setup facilitates efficient inter-server communication with a bandwidth capacity of 800Gbps.
Software.
Experiments are implemented in PyTorch 2.1.1 and Triton 2.0.0 with CUDA 11.7, cuDNN 8.0, and NCCL 2.14.3. Our algorithm is developed upon Metaseq and DeepSpeed.
A.3 Experimental Setup
@@
The training configuration is set with specific hyperparameters: a learning rate of 0.0005 to control the optimization step size, a cap of 50,000 updates to define the training duration, and a 2,000-update warmup period to stabilize early training by gradually adjusting the learning rate. Additionally, a weight decay rate of 0.01 is used for regularization to avoid over-fitting (Sun et al., 2024). The Adam optimizer, with beta values of 0.9 and 0.999, is chosen for managing the momentum and scaling of gradients, aiding in effective and stable training convergence (Zhou et al., 2020). Different DDP backends, including PyTorch DDP (abbr. DDP), Legacy DDP, FSDP, ZeRO-series, are selected in experiments for cross-validation of compatibility with LASP.
A.4 Generalization of LASP
While LASP is initially inspired by linear attention mechanisms, we aim to show its broader applicability to various linear sequence modeling approaches. This section investigates the generalization of LASP through both theoretical analysis and empirical validation.
In the theoretical aspect, we first define the following terms: Memory State $\mathbf{m}_{t}\in\mathbb{R}^{k\times d}$, Input State $\mathbf{i}_{t}\in\mathbb{R}^{d}$, Expand State $\mathbf{e}_{t}\in\mathbb{R}^{k}$, Oscillation State $\mathbf{o}_{t}\in\mathbb{R}^{k\times m}$, Shrink State $\mathbf{s}_{t}\in\mathbb{R}^{k}$ and write a general form of recurrent memory as (Qin et al., 2024b)
$\mathbf{m}_{t}=\mathbf{o}_{t}{\mathbf{m}}_{t-1}+{\mathbf{e}}_{t}{\mathbf{i}}_{t}^{\top}.$ (24)
which is general form of the recurrence form of Linear Attention in Eq. (5) with specified $\mathbf{o}_{t}$ and $\mathbf{e}_{t}$:
$\mathbf{kv}_{t}=\mathbf{\lambda}\mathbf{kv}_{t-1}+\mathbf{k}_{t}\mathbf{v}_{t}^{\top}.$ (25)
The design of LASP can be seamlessly applied to models which is able to be generally expressed by Eq. (24). These models include: S4 (Gu et al., 2022), S5 (Smith et al., 2022), DSS (Gupta et al., 2022), TNN (Qin et al., 2023a), Linear Attention (Katharopoulos et al., 2020), TNL (Qin et al., 2024a), RetNet (Sun et al., 2023), Mamba (Gu & Dao, 2023), RWKV-4 (Peng et al., 2023), Cosformer (Qin et al., 2022b), Lrpe (Qin et al., 2023c), GLA (Yang et al., 2023), GateLoop (Katsch, 2023), DUR (Mao, 2022), GFW (Schlag & Schmidhuber, 2018), HGRN (Qin et al., 2024f; e), and LRN (Martin & Cundy, 2018). We list all these models and their corresponding elements in Table 5.
Table 5: Checklist for Typical Linear-Complexity Sequence Modeling Methods within the Defined General Form. For each method, the following states are outlined: Input State, Expand State, Oscillation State, Shrink State, and Memory State. If the state is directly linked to the input sequence, the subscript i is emphasized. Note that we use $\mathbf{1}^{(k)}\in\mathbb{R}^{k}$, where $\mathbf{1}^{(k)}_{j}=1$ for $j=1,\ldots,k$, and $\mathbf{J}^{(kd)}=\mathbf{1}^{(k)}{\mathbf{1}^{(d)}}^{\top}\in\mathbb{R}^{k\times d}$.
Method Input $\mathbf{i}_{t}$ Expand $\mathbf{e}_{t}$ Oscillation $\mathbf{o}_{t}$ Shrink $\mathbf{s}_{t}$ Memory $k\times d$
S4 $\mathbf{x}_{t}$ $\mathbf{B}$ $\mathbf{A}$ $\mathbf{C}$ $k\times 1$
S5 $\mathbf{x}_{t}$ $\mathbf{B}$ $\mathbf{A}$ $\mathbf{C}$ $k\times d$
DSS $\mathbf{x}_{t}$ $\mathbf{B}$ $\mathbf{a}\mathbf{1}_{k}^{\top}$ $\mathbf{C}$ $k\times d$
TNN $\mathbf{x}_{t}$ $\mathbf{B}$ $\mathbf{A}$ $\mathbf{C}$ $k\times d$
Linear Attention $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\mathbf{J}^{(kd)}$ $\mathbf{C}_{t}$ $k\times d$
TNL/RetNet $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\lambda\mathbf{J}^{(k)}$ $\mathbf{C}_{t}$ $k\times d$
Mamba $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\mathbf{A}_{t}$ $\mathbf{C}_{t}$ $k\times d$
RWKV4 $\mathbf{x}_{t}$ $\exp(\mathbf{k}_{t})$ $\exp(-w)$ $\mathbf{C}_{t}$ $1\times 1$
Cosformer $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\exp(i\theta)\mathbf{J}^{(kd)}$ $\mathbf{C}_{t}$ $k\times d$
LRPE $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\exp(i\Theta){\mathbf{1}^{(d)}}^{\top}$ $\mathbf{C}_{t}$ $k\times d$
GLA/GateLoop $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\mathbf{g}_{t}\mathbf{1}_{d}^{\top}$ $\mathbf{C}_{t}$ $k\times d$
DUR/GFW $\mathbf{x}_{t}$ $\mathbf{B}_{t}$ $\mathbf{g}_{t}\bar{\mathbf{g}}_{t}^{\top}$ $\mathbf{C}_{t}$ $k\times d$
HGRN/LRN $\mathbf{x}_{t}$ $1-\mathbf{A}_{t}$ $\mathbf{A}_{t}$ $\mathbf{C}_{t}$ $1\times 1$
We also give the complete explanation for each modeling method as below.
S4. In S4, we obtain $u_{t}\in\mathbb{R}^{d}$ through linear projection from input $x_{t}$ and $A\in\mathbb{R}^{k\times k},B,C\in\mathbb{R}^{k\times 1}$ through SSM parameterization. The calculation is as follows:
$m_{t}=Am_{t-1}+Bu_{t}^{\top},\ y_{t}=m_{t}^{\top}C.$
Note that the original definition of S4 is defined as a channel-wise mappings $f_{i},i=1,\ldots,d$ of $\mathbb{R}^{n\times 1}\to\mathbb{R}^{n\times 1}$.
S5. The recurrence equation of S5 is the same as S4, with the only difference being the direct definition of the mapping $\mathbb{R}^{n\times d}\to\mathbb{R}^{n\times d}$ and $B,C\in\mathbb{R}^{k\times d}$.
DSS. The recurrence equation of DSS is same as S4/S5, with the only difference being the direct definition of the mapping $\mathbb{R}^{n\times d}\to\mathbb{R}^{n\times d}$ and $B,C\in\mathbb{R}^{k\times d},A=\mathrm{Diag}{a}\in R^{k\times k}$.
TNN. According to (Qin & Zhong, 2023), TNN can be losslessly converted to SSM, where $C=J^{(kd)}\in\mathbb{R}^{k\times d},B\in\mathbb{R}^{k\times d},A=\mathrm{Diag}{\lambda_{1},\ldots,\lambda_{k}}\in\mathbb{R}^{k\times k}$, get $u_{t}$ from $x_{t}$ through linear projection, and it can be expressed as a recursive formula:
$m_{t}=Am_{t-1}+Bu_{t}^{\top},\ y_{t}=m_{t}^{\top}C.$
Linear Attention. In Linear Attention, we obtain query $q_{t}\in\mathbb{R}^{k}$, key $k_{t}\in\mathbb{R}^{k}$, value $v_{t}\in\mathbb{R}^{d}$ from the input $x_{t}\in\mathbb{R}^{d}$ through linear projection, and recursively calculation is as follows:
$kv_{t}=kv_{t-1}+k_{t}v_{t}^{\top},\ y_{t}=kv_{t}^{\top}q_{t}.$
TNL/RetNet. TNL/RetNet is a form of Linear Attention with exponential decay and the method for getting $q_{t},k_{t},v_{t}$ are the same as those in Linear Attention, and lambda is a predefined parameter that cannot be learned. Its recursive calculation is:
$kv_{t}=\lambda kv_{t-1}+k_{t}v_{t}^{\top},\ y_{t}=kv_{t}^{\top}q_{t}.$
Mamba. Mamba can be seen as a data-dependent S4. It uses the similar method to get $u_{t},A,B,C$, the ${A_{t}},B_{t},C_{t}$ are computed throuth $x_{t}$ and $A,B,C$. Its recurrence equation is defined as:
$m_{t}=A_{t}\odot m_{t-1}+B_{t}u_{t}^{\top},\ y_{t}=m_{t}^{\top}C_{t}.$
RWKV-4. In RWKV-4, we get $r_{t},k_{t},v_{t}$ through linear projection from input $x_{t}$ and $w$ as a learnable weight. Ignoring the denominator of RWKV-4, the recurrence equation can be simplified as:
$m_{t}=\exp(-w)m_{t-1}+\exp(k_{t})v_{t}^{\top},\ y_{t}=m_{t}^{\top}r_{t}.$
Similar to S4, RWKV4 uses channel-wise mapping $f_{i},i=1,\ldots,d$ of $\mathbb{R}^{n\times 1}\to\mathbb{R}^{n\times 1}$.
Cosformer. In Cosformer, we obtain query $q_{t}\in\mathbb{R}^{k}$, key $k_{t}\in\mathbb{R}^{k}$, value $v_{t}\in\mathbb{R}^{d}$ from the input $x_{t}\in\mathbb{R}^{d}$ and a prefined $\theta$(not learnable). Then recursively calculate as follows:
$kv_{t}=\exp(i\theta)kv_{t-1}+k_{t}v_{t}^{\top},\ y_{t}={Rel}[{kv}_{t}^{\top}]q_{t}.$
Lrpe. In Lrpe, we obtain query $q_{t}\in\mathbb{R}^{k}$, key $k_{t}\in\mathbb{R}^{k}$, value $v_{t}\in\mathbb{R}^{d}$ from the input $x_{t}\in\mathbb{R}^{d}$, $\theta$ as a learnable weight and recursively calculate as follows:
$kv_{t}=\Lambda kv_{t-1}+k_{t}v_{t}^{\top},\Lambda=\mathrm{diag}(\exp(i\theta_{1}),\ldots,\exp(i\theta_{k})),y_{t}=\mathrm{Rel}{[{kv}]_{t}}^{\top}q_{t}.$
GLA/GateLoop. In GLA/GateLoop, we obtain query $q_{t}\in\mathbb{R}^{k}$, key $k_{t}\in\mathbb{R}^{k}$, value $v_{t}\in\mathbb{R}^{d}$, decay $g_{t}\in R^{k}$ from the input $x_{t}\in\mathbb{R}^{d}$ and recursively calculate as follows:
$kv_{t}=\mathrm{Diag}(g_{t})kv_{t-1}+k_{t}v_{t}^{\top},y_{t}=kv_{t}^{\top}q_{t}.$
DUR/GFW In DUR/GFW, we obtain query $q_{t}\in\mathbb{R}^{k}$, key $k_{t}\in\mathbb{R}^{k}$, value $v_{t}\in\mathbb{R}^{d}$, decay $g_{t}\in R^{k},\bar{g}_{t}\in\mathbb{R}^{d}$ from the input $x_{t}\in\mathbb{R}^{d}$, and recursively calculate as follows:
$kv_{t}=(g_{t}\bar{g}_{t}\top)\odot kv_{t-1}+k_{t}v_{t}^{\top},y_{t}=[kv]_{t}^{\top}q_{t}.$
HGRN/LRN In HGRN/LRN, we obtain output gate $o_{t}\in\mathbb{R}^{1}$, forget gate $f_{t}\in\mathbb{R}^{1}$, input state $i_{t}\in\mathbb{R}^{1}$ from the input $x_{t}\in\mathbb{R}^{1}$, and recursively calculate as follows:
$h_{t}=f_{t}\odot h_{t-1}+(1-f_{t})i_{t}^{\top},y_{t}=h_{t}^{\top}o_{t}.$
Similar to S4, HGRN/LRN use channel-wise mapping $f_{i},i=1,\ldots,d$ of $\mathbb{R}^{n\times 1}\to\mathbb{R}^{n\times 1}$.
To empirically validate the generalization of LASP, we adopt the experimental setup from Table 2 and apply LASP to three additional linear sequence modeling methods listed in Table 5, namely Cosformer, RetNet, and Mamba. The convergence results, presented in Table 6, demonstrate that LASP does not negatively affect convergence and achieves performance on par with the baselines.
Table 6: Convergence Results of LASP on Cosformer, RetNet and Mamba. TNL with 0.4B parameters are tested with a batch size of 2 and sequence length of 16K.